# Precise Interception Flight Targets by Image-based Visual Servoing of Multicopter

### v1

@@ IV-A Platform introduction and experimental design
In this section, the proposed algorithm is validated and compared using SITL simulation experiments with RflySim [30].These simulations replicate the architecture and algorithms of real flight experiments, ensuring high fidelity and ease of portability. First, a multi-angle static target interception experiment is designed to verify the accuracy of the proposed algorithm. Next, moving target interception experiments are performed and compared with other algorithms. These experiments include targets with three common motion models and initial relative positions in three different directions.
@@ IV-B Static Target Intercept Simulation Experiments
Fifty static target interception experiments are performed with targets at various locations. The multicopter starts at the position (0,0,10) meters in the EFCS, while the target positions, as shown in Fig. 5(a), are randomly generated between 15 and 35 meters from the multicopter. The multicopter’s initial velocity is set to be directly forward, and $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=-0.03$, $k_{p}=-0.01$ are set.
Experimental results demonstrate that the proposed algorithm accurately intercepts most targets within the FOV. The target distribution in the image plane before interception is shown in Fig. 5(c). The interception error box and CEP are illustrated in Fig. 5(b) and Fig. 5(d), respectively. Compared to previous works [18] and [19], which reported $\mathrm{CEP}_{\left[18\right]}=0.457$ m and $\mathrm{CEP}_{\left[19\right]}=0.332$ m, the proposed algorithm achieves a $\mathrm{CEP}_{\mathrm{Proposed}}=0.089$ m, representing reductions of 80.5% and 73.2%, respectively. This indicates that the proposed algorithm enhances the accuracy of static target interception from the decimeter level to the centimeter level in the simulation experiments.
@@ IV-C Moving Target Intercept Simulation Experiments
This experiment evaluates the interception accuracy of the proposed algorithm across various target maneuvering scenarios. Three maneuvering models were tested: constant velocity (CV), constant acceleration (CA), and sinusoidal maneuver (SM). Parameters were consistent with those used for intercepting stationary targets. The initial relative positions included three interception scenarios: top-to-bottom, flanking, and tailgate, as shown in Fig. 6. Initially, the interceptor multicopter hovers at (0, 0, 10) m. The target follows three different maneuvering models starting from (0, 25, 1) m, (-8, 15, 8) m, and (0, 30, 10) m, respectively. The models are expressed as:
@@
Experimental results demonstrate that the proposed algorithm achieves the highest interception accuracy in the simulations. Fig. 7 show the variations in normal acceleration and velocity during interception. The proposed algorithm maintains low normal acceleration and velocity, which is ideal for minimizing interception errors. Low normal acceleration reduces trajectory deviations, while low normal velocity minimizes the impact of image processing delays on accuracy. Interception errors are illustrated in Fig. 8. Table I confirms that the proposed algorithm outperforms the other two algorithms.
@@ V-A Introduction to Flight Experiments
Real flight experiments are designed to verify the robustness and effectiveness of the proposed algorithm under more complex conditions. The algorithms, architectures, and parameter settings are the same as in the simulation: $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=-0.03$, $k_{p}=-0.01$. The real-world experiments are depicted in Fig. 10. The intruder is a red balloon attached to an multicopter by a rope. The interceptor is another multicopter equipped with several components. At the front is a Jetson CSI monocular camera with a 120-degree FOV for capturing images. Behind it is an NVIDIA Jetson Xavier NX, which serves as the onboard computer for processing sensor information and generating control commands. A Pixhawk Nano V5 handles flight control, generating PWM signals to the motors. Additionally, the interceptor is equipped with an RTK module for data analysis target localization (note: this module is not used during the interception process). A video of the flight experiments can be found at https://youtu.be/sQdRCgnRp4g.
@@
V-B Static Target Intercept Flight Experiment
The interceptor drone starts from a standstill to intercept a red balloon shaking in the wind about 10m away. The results of the experiment are shown in Figure 11. And the trajectory is shown in (a) and (b). As shown in (c), the target always stays in the FOV during interception and moves only a small range. It proves the effectiveness of the proposed FOV holding controller in the real world. The normal acceleration change little during the whole interception process can be seen in (d). And As shown in (e), normal velocity keep at a small level and the maximum velocity is more than 6m/s.
@@ V-B Static Target Intercept Flight Experiment
V-C Moving Target Intercept Flight Experiments
Three interception experiments of different maneuver targets are designed: accelerated escape, uniform takeoff, and circular trajectory. These experiments are challenging due to the manoeuvrability of the target and the effects of wind in the field. The relevant initial settings are the same as for the static interception experiment. The interceptor multicopter start the interception task from a standstill. Scenarios and interception trajectory can be seen in Fig. 12. The starting positions are about 10m according to the target.
Figure 12: Schematic of (a) accelerated escape, (b) uniform takeoff, and (c) circular trajectory flight tagets and interceptor trajectories .
All three intercept experiments successfully achieved interception of maneuvering targets. The time taken for interception was approximately 4.8s, 3.9s, and 4.6s, respectively. The velocity and normal acceleration during interception can be seen in Fig. 13. At the end stage of interception, the normal acceleration and normal velocity of the interceptor were essentially the same as those in the hovering stage. The final interception speeds were around 5 m/s. During the interception of the accelerated escape target and the circular trajectory target, a significant portion of the interceptor’s velocity was used for normal velocity in the early stage. This indicates that the PNG works to ensure a flat trajectory by constantly adjusting the relative position before interception. Recorded trajectories further demonstrate the effectiveness of the proposed guidance law in intercepting maneuvering targets.
@@ V-C Moving Target Intercept Flight Experiments
Additionally, all experiments maintained the target within the FOV during interception. The position of the target in the image plane can be seen in Fig. 14. Generally, all targets moved within a small area in the middle of the image. Despite some dispersion in the first experiment when intercepting the accelerated escape target, the results prove the effectiveness and robustness of the proposed FOV holding controller in intercepting maneuvering targets in real scenarios.
@@
V-D Compare Intercept Flight Experiments
However, the two most relevant algorithms [19] and [22] were tested for intercepting an accelerated escape target. This maneuver model is considered the most challenging by previous experiments. As shown in Fig. 15, both algorithms failed during the interception and were subsequently returned by manual remote control. After analysis, the reasons for the interception failure of these two algorithms may be: (1) The overload during interception exceeds the dynamics range of the multicopter, causing the target to exceed the FOV. (2) Random wind disturbances during the experiments result in irregular balloon oscillations. (3) Two algorithms’ interception accuracy are larger than the size of the balloon.
@@ VI Conclusion
In this paper, an IBVS algorithm based on PNG was designed to address the problem of intercepting non-cooperative flight targets with a multicopter. The flat trajectory generated by PNG reduces errors due to image processing latency, thereby improving interception accuracy. Additionally, the proposed algorithm demonstrates increased robustness through the FOV holding controller designed for the multicopter. Extensive simulation and real flight experiments have validated the accuracy of the proposed algorithm’s interception capability. However, the proposed algorithm may face challenges when intercepting higher-speed targets. Therefore, future work will focus on enhancing guidance law applications and employing more precise control methods.

### v2

@@ IV-A Platform Introduction and Experimental Design
In this section, the proposed algorithm is validated and compared using SITL simulation experiments with RflySim [39].These simulations replicate the architecture and algorithms of real flight experiments, ensuring high fidelity and ease of portability. Initially, a multi-directional static target interception experiment is designed to verify the accuracy of the proposed algorithm. Next, moving target interception experiments are performed and compared with other algorithms. These experiments include targets with three common motion models and initial relative positions in three different directions.
@@ IV-B Static Target Intercept Simulation Experiments
Fifty static target interception experiments are performed with targets at various locations. The multicopter starts at the position (0,0,10) meters in the EFCS, while the target positions, as shown in Fig. 5(a), are randomly generated between 15 and 35 meters from the multicopter. The multicopter’s initial velocity is set to move directly forward, and $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=0.03$, $k_{d}=0.01$ are set.
The recorded data from the proposed controller during static target interception are shown in Fig. 5(a)–(d). The distributions of 50 targets in 3D space and the image plane are shown in (a) and (c), respectively. Interception accuracy is quantified by the Circular Error Probability (CEP), which indicates the radius within which the interceptor has a 50% probability of hitting the target [40]. The CEP is $\mathrm{CEP}_{\mathrm{Proposed}}=0.089$ m, as shown in (b) and (d). This represents a substantial improvement over previous works [26] and [28], reporting CEP values of $\mathrm{CEP}_{\left[26\right]}=0.457$ m and $\mathrm{CEP}_{\left[28\right]}=0.332$ m, respectively. The proposed method achieves reductions of 80.5% and 73.2%, demonstrating a significant improvement in accuracy. Notably, this result marks a transition from decimeter-level performance to centimeter-level accuracy in this classic test scenario.
@@ IV-C Moving Target Intercept Simulation Experiments
This experiment evaluates the interception accuracy of the proposed algorithm across various target maneuvering scenarios. Three maneuvering models were tested: constant velocity (CV), constant acceleration (CA), and sinusoidal maneuver (SM), each modeling different types of evasive actions typically encountered in low-altitude target interception tasks. Parameters were consistent with those used for intercepting stationary targets. The initial relative positions included three interception scenarios: top-to-bottom, flanking, and tailgate, as shown in Fig. 6. These scenarios test the algorithm’s performance in diverse dynamic environments, challenging the controller’s robustness in both direct pursuit and evasive maneuvers. Initially, the interceptor multicopter hovers at (0, 0, 10) m. The target follows three different maneuvering models starting from (0, 25, 1) m, (-8, 15, 8) m, and (0, 30, 10) m, respectively. The models are expressed as:
@@
As shown in Fig. 6, the proposed algorithm achieves smoother trajectories when intercepting moving targets. The lower normal acceleration and velocity in the final phase of interception, shown in Fig. 7, demonstrate the effectiveness of PNG in the proposed IBVS controller. Interception errors are illustrated in Fig. 8. Although the trajectory smoothness and normal velocity of PG are lower than those of the proposed method in some scenarios, PG experiences a surge in normal acceleration towards the end, leading to reduced interception accuracy, as shown in Fig. 7 and Fig. 8. Table I confirms that the proposed algorithm’s interception accuracy outperforms the other two algorithms in intercepting moving targets.
@@ V-A Introduction to Flight Experiments
Extensive SITL experiments demonstrate the interception accuracy and effectiveness of the proposed algorithm. In this section, a series of experiments will be conducted to verify the effectiveness and robustness of the proposed algorithm in different real scenarios.
In all experiments, the parameter settings were kept consistent with the simulation: $k_{a}=2$, $K_{y}=K_{z}=3$, $k_{p}=0.03$, $k_{d}=0.01$, and the red balloon was used instead of the target. The indoor flight experiments under different wind speeds were designed by first considering the potential effects of outdoor wind conditions on both the target balloon and the vehicle. Following this, the experimental scenarios were set outdoors. Static target interception, moving target interception, and comparison experiments were conducted.
V-B Indoor Experiments with Different Wind Conditions
The indoor experiments were conducted using the experimental platform shown in Fig. 10. The multicopter is equipped with a Jetson CSI monocular camera ($120^{\circ}$ FOV) for capturing images, a NVIDIA Jetson Xavier NX for processing sensor data, and a Pixhawk 6C Mini for flight control. The fans can provide winds of 2 to 6 m/s from different distances, and the target balloon is attached to the base by a soft rope to swing with the wind. The target balloon had a diameter of about 16 cm, and the multi-rotor drone had a wheelbase of 15 cm, making it a very challenging series of tests. A total of 20 interception experiments were conducted at different wind speeds. The video of the experiments can be found at https://youtu.be/vl2UTgEjeyk or https://b23.tv/cx5egf4.
Figure 10: Indoor interception experiment scenarios under different wind conditions.
To ensure experimental consistency, the same balloon was reused across all tests, with the Optitrack motion capture system providing accurate initial position data for the interceptor. The multicopter began at the position (0, 6, 1.5) m and intercepted the balloon located at (0.5, 2.3, 1) m. Wind speed at the balloon’s location was measured using an anemometer when the balloon was stationary. A series of five consecutive experiments were conducted at wind speeds of 0 m/s, 2 m/s, 4 m/s, and 6 m/s to evaluate the robustness of the controller and the interception success rate.
The median, quartiles, boundaries, and outliers of the target in the image plane are shown in Fig. 11, while the Interquartile Range (IQR, defined as the difference between the upper and lower quartiles) and the interception success rate are presented in Table II. As the wind speed increases, both the median of $u$ and the IQR increase, indicating larger errors in the yaw angular rate controller. Conversely, the median and IQR of $v$ initially increase, then decrease (or vice versa), suggesting that the wind initially caused the target balloon to float upward, bringing it closer to the center of the image plane. As the wind speed continued to rise, the balloon gradually deviated from the image center. When the wind speed was below 4 m/s, the proposed controller achieved an interception success rate of over 80%. However, at a wind speed of 6 m/s, the success rate dropped to 40%.
Overall, although the interception success rate and robustness of the proposed system decrease as wind speed increases, it still demonstrates effectiveness in breezy conditions. Moreover, unlike balloons, real-world non-cooperative targets (e.g., drones, birds, kites, etc.) do not oscillate with the wind in the same manner.
Figure 11: Comparison of normalized planar image errors at different wind speeds.
TABLE II: Normalized image plane distribution data and interception success rate under different wind conditions.
Scenarios median of ($u$,$v$) IQR of ($u$,$v$) Success rate
0 m/s (-0.002,-0.064) (0.019,0.058) 100%
2 m/s (-0.010,-0.063) (0.046,0.043) 80%
4 m/s (-0.014,-0.060) (0.061,0.057) 80%
6 m/s (-0.025,-0.075) (0.145,0.075) 40%
V-C Outdoor Experiments with Different Targets
To thoroughly evaluate the performance of the proposed controller in real-world scenarios, we designed experiments to intercept both static targets and three types of maneuvering targets in the field. The three types of maneuvering targets are designed as accelerated escape, uniform takeoff, and circular trajectory. The wind speed during the experiment ranged from 0 to 5 m/s.
The outdoor experiments are shown in Fig. 12. The intruder is a red balloon, similar to the one used in the indoor experiments, attached to a multicopter by a rope. The interceptor is another multicopter, equipped with various sensors and actuators. The hardware architecture of the interceptor is similar to that of the indoor experiment, featuring a larger wheelbase and a higher thrust-to-weight ratio. The multicopter has a thrust-to-weight ratio of approximately 3 and is capable of providing a maximum acceleration of 5 $\mathrm{m/s^{2}}$ to successfully intercept the target. Additionally, we oriented the camera horizontally, which increased the multicopter’s field of view in the vertical direction, as shown in Fig. 16. The interceptor multicopter start the interception task from a standstill from about 15 m away. A video of the flight experiments can be found at https://youtu.be/sQdRCgnRp4g or https://b23.tv/nBPFA7j.
@@ V-C Outdoor Experiments with Different Targets
In the first flight experiment to intercept a static target, the interceptor successfully intercepted the target after approximately 6.2 s, reaching a terminal speed of 6.3 m/s. The results of the experiment are shown in Fig. 13. Due to the field winds, the balloon experienced slight movement, requiring continuous velocity direction adjustment by the interceptor’s controller, as shown in the trajectories in (a) and (b). Additionally, due to the wind, a normal acceleration of approximately 2 m/s is required to achieve hovering, as shown in (d) at $t=0$ s. (d) and (e) show that the normal acceleration increments are small and the normal velocity remains low throughout the interception process, demonstrating the effectiveness of the proposed PNG-based controller. The effectiveness of the proposed FOV holding controller is also demonstrated by the fact that the target moved only a small distance in the image plane, as shown in (c).
In the next three flight experiments intercepting maneuvering targets, the interceptor successfully intercepted the target along smooth trajectories, as shown in Fig. 14. Fig. 15 show the velocity and normal acceleration during the interception process. During the interception of the accelerated escape target and the circular trajectory target, a significant portion of the interceptor’s velocity was used for normal velocity in the early stage due to drastic changes in line-of-sight angle caused by large lateral movement of the target. In addition, the normal acceleration of interceptor at the end phase was essentially the same as that in the hovering stage. These results indicate the PNG works to ensure interception accuracy. All three scenarios achieve effective interception at a terminal velocity of about 5 m/s. Additionally, all three experiments kept the target within the field of view (FOV) throughout the interception process, as shown in Fig. 16. Overall, all targets remained within a small region at the center of the image. Despite some deviation in the first experiment when intercepting the accelerated escape target, the results demonstrate the effectiveness and robustness of the proposed FOV holding controller in intercepting maneuvering targets in real scenarios.
Figure 14: Schematic of (a) accelerated escape, (b) uniform takeoff, and (c) circular trajectory flight targets and interceptor trajectories.
@@
However, the two most relevant and comparable algorithms [28] and [27] were tested on intercepting an accelerated escape target. This maneuver model has been regarded as the most challenging by previous studies. As shown in Fig. 17, both algorithms failed to complete the interception successfully and were manually returned by remote control after failure. Upon analysis, the reasons for the interception failure of these two algorithms are identified as: (1) The overload during interception exceeds the dynamic range of the multicopter, causing the target to move out of the FOV. (2) Random wind disturbances during the experiments cause irregular balloon oscillations. (3) The interception accuracy of the two algorithms exceeded the size of the balloon.
@@ VI Conclusion
This paper presents an advanced Image-Based Visual Servoing (IBVS) algorithm integrated with proportional navigation guidance (PNG) for the precise interception of non-cooperative aerial targets using a multicopter. The proposed algorithm leverages the flat trajectory generated by PNG, effectively mitigating errors caused by image processing latency and significantly enhancing interception accuracy. Furthermore, the incorporation of a Field of View (FOV) holding controller ensures continuous and stable target tracking, even in dynamic and unpredictable environmental conditions. Through extensive simulations and real-flight experiments under various scenarios, the proposed algorithm has demonstrated its superior performance in terms of accuracy and robustness when compared to existing methods. However, the algorithm’s performance may degrade when intercepting high-speed targets due to challenges in maintaining trajectory stability at high velocities. In future work, we plan to refine the guidance law by incorporating adaptive strategies, real-time state estimation techniques, and more precise control methods to further improve the interception performance, particularly in complex and high-speed target interception scenarios.